When a constant force acts upon an object, the acceleration of the object varies inversely with its mass. When a certain constant force acts upon an object with a massive 26 kg the acceleration of the object is 3 m/s^2. This is the same force acts upon another object with mass is 39 kg what is the objects acceleration.
Answer #1

F/m = a or F = ma

For the first statement, plugging in the mass and acceleration gives:

F = 26 x 3 = 78 Newtons (where Newtons is the standard unit of force)

Applying the force above to the second mass yields:

a = 78/39 = 2m/s^2.

Please rate my answer with thumbs up if it helps!